# Think-at-Hard: Selective Latent Iterations to Improve Reasoning Language Models

### v1

@@ 4.2 Model Architecture
During inference, token $i$ verbalizes when $c_{i}^{(d)}$ falls below threshold $c_{\text{threshold}}$ or reaches maximum depth $d_{\max}$.
@@ 4.3 Training Scheme
We employ a two-stage training scheme: first finetune the backbone model for dynamic iteration, then the iteration decider, all using an oracle policy.
Motivation. As shown in Figure 2(c), the backbone network $\theta_{d}$ and the neural iteration decider $\mathcal{I}_{\phi}$ are tightly coupled: the backbone generates hidden states as inputs for the decider, while the decider controls the backbone’s KV cache and iterations. Training both simultaneously causes instability due to mutual distribution shifts. Therefore, we adopt a stable two-stage approach where both components are sequentially trained to align an oracle iteration policy $\pi$.
Oracle Iteration Policy $\pi$. To guide training, we define an oracle policy $\pi$ that determines token difficulty using a frozen reference LLM, following fu2025r2r. A token is classified as easy if the reference model correctly predicts it with a single forward pass, and hard otherwise. Throughout the paper, we use the supervised fine-tuned (SFT) variant of the base model as the reference model.
Formally, let $\hat{t}_{i+1}$ denote the reference model’s top-1 prediction and $t_{i+1}$ the ground-truth token. For explanation simplicity, we assume maximum iteration depth $d_{\max}=2$ in Equation 7; the general case is detailed in Appendix A.2.1. The oracle iteration depth $d^{\pi}$ is:
$d_{i}^{\pi}\;=\;1+\bm{1}\!\left[\hat{t}_{i+1}\neq t_{i+1}\right],$ (7)
where $\bm{1}[\cdot]$ is the indicator function. The per-depth continuation label becomes:
$c_{i}^{(d)}\;=\;\bm{1}\!\left[d\leq d_{i}^{\pi}\right],$ (8)
indicating whether iteration should continue at depth $d$. Table 4 and Figure 1 verify the effectiveness of the oracle policy.
@@
We freeze the backbone model $(\theta,\Delta)$ and train the iteration decider $\phi$ to imitate the oracle policy’s continuation decisions. We minimize binary cross-entropy with class reweighting for label imbalance:
$\mathcal{L}_{\text{dec}}(\phi)\;=\;-\sum_{i}\sum_{d=1}^{\min\{d_{\max}-1,d_{i}^{\pi}\}}\Big[w^{(d)}_{\text{stop}/\text{cont.}}c_{i}^{(d)}\log\hat{c}_{i}^{(d)}+\big(1-c_{i}^{(d)}\big)\log\big(1-\hat{c}_{i}^{(d)}\big)\Big],$
where $c_{i}^{(d)}$ is the ground-truth continuation label, $\hat{c}_{i}^{(d)}$ is the predicted probability, and $w^{(d)}_{\text{stop}/\text{cont.}}$ is the occurrence ratio of stop label divided by continue label, respectively.
Our two-stage scheme stabilizes training by decoupling backbone learning (conditioned on a fixed $\pi$) from policy learning (imitation of $\pi$).
@@ 5.1 Setup
We present key configurations here, with more detailed setups in the Appendix.
Baselines. We compare diverse methods under equal parameter budgets, using Qwen3-0.6B-Base and Qwen3-1.7B-Base (yang2025qwen3) as backbones. We compare TaH over several fixed-depth strategies: (1) Standard, which always verbalizes directly and reduces to the original Qwen model; (2) AlwaysThink, which applies the maximum number of latent iterations to all tokens; (3) SoftThink, following official baseline implementation (zhang2025soft) on top of the Standard model. Unless otherwise specified, both TaH and AlwaysThink use a maximum of two iterations. We also compare with dynamic query routing via matrix factorization (ong2024routellm), routing between MobileLLM-R1-360M (zhao2025mobilellm) and Qwen3-1.7B, as well as between Qwen3-0.6B and Qwen3-4B, to match average parameter sizes of 0.6B and 1.7B.
TaH Setup. Before training, we prune one layer from the base model so that TaH matches the parameter count of baselines. The layer is chosen to minimize the increase in validation loss. We also report results for an unpruned variant, TaH+, which adds less than 3% extra parameters from LoRA and iteration decider. The detailed parameter composition is shown in Table 11. Following (fu2025r2r), we set the continuation threshold $c_{\text{threshold}}=0.9$ with about 6% of tokens being iterated twice. The oracle policy $\pi$ uses Qwen3-0.6B and 1.7B as reference models to determine token difficulty during training.
Training Scheme. All models are trained on the math subset of Open-R1 (openr1) using supervised finetuning. To fit memory and compute limits, we exclude responses longer than 8,192 tokens; all other training settings follow the official Open-R1 script. The filtered dataset contains 300M tokens, with 1% reserved for validation. Each method is sufficiently trained for 5 epochs, and we select the checkpoint with the lowest validation loss as the final model. All backbones are initialized from the corresponding Qwen3-Base.
Evaluation Setup. We evaluate across challenging reasoning benchmarks, including GSM8K (cobbe2021training), MATH500 (hendrycks2021measuring), AMC23 (American Mathematics Competitions), AIME25 (American Invitational Mathematics Examination), and OlympiadBench (he2024olympiadbench). The maximum generation length is set to 8,192 tokens for all benchmarks, except GSM8K which uses 4,096 due to its simpler problems and larger size. Performance is reported as pass@1 under a zero-shot chain-of-thought setting, using sampling temperature 0.6. For large datasets (MATH500, OlympiadBench, GSM8K), we generate one sample per problem; for small datasets (AMC23, AIME25), we generate eight samples per problem.
@@ 5.2 Performance
Table 1: Accuracy comparison of different baselines across five benchmarks and two model sizes. Subscripts indicate improvement over Standard.
Method
Param. Benchmark Standard Routing SoftThink AlwaysThink TaH TaH+
0.6B GSM8K 62.5 45.6 61.3 54.6 64.4 68.8
MATH500 47.2 27.3 48.8 32.8 51.2 54.2
AMC23 23.4 10.9 24.1 20.3 32.5 30.6
AIME25 4.2 1.0 2.5 1.5 4.2 5.0
OlympiadBench 18.8 7.4 19.4 10.2 23.9 24.0
Average 31.2 18.5 31.2 23.9 35.2/+/+4.0 36.5/+/+5.3
1.7B GSM8K 82.1 71.2 79.6 79.3 84.5 85.8
MATH500 68.4 60.0 68.8 61.8 74.4 73.0
AMC23 42.2 42.2 43.1 42.5 48.4 51.2
AIME25 13.3 10.2 12.9 10.0 17.9 14.6
OlympiadBench 33.0 30.6 33.4 30.0 38.8 41.2
Average 47.8 36.8 47.6 44.7 52.8/+/+5.0 53.2/+/+5.4
Benchmark Evaluation. We validate TaH’s reasoning ability through extensive tests across five challenging math benchmarks. Table 1 presents performance results for models at 0.6B and 1.7B parameter sizes. Starting from strong Qwen3-Base models, we observe that existing approaches show limited effectiveness: fixed-depth recurrent transformers (AlwaysThink) and query routing fail to consistently outperform the standard direct-answer baseline. SoftThink provides improvements on some cases, yet remain marginal overall. In contrast, TaH achieves consistent gains, delivering average improvements of 4.0% and 5.0% for the 0.6B and 1.7B models, respectively. Our enhanced variant (TaH+), which only adds less than 3% additional parameters, pushes these gains to 5.3% and 5.4%. Relative to AlwaysThink, the gains are 8.1-11.3% for TaH, and 8.5-12.6% for TaH+.
Figure 3: Training dynamics of the LLM backbone on Qwen3-0.6B-Base. TaH converges rapidly and achieves lower perplexity.
Figure 4: GSM8K accuracy with respect to continuation threshold. Numbers in brackets indicate the percentage of tokens that iterate twice.
Training Dynamics. During stage 1 (LLM backbone training), guided by the oracle policy that only triggers a second iteration on hard tokens, TaH converges notably faster than the Standard baseline. It also achieves much lower perplexity on the validation dataset as shown in Figure 3. During stage 2 (iteration-decider training), the neural decider successfully imitates the oracle strategy, reaching about 83% accuracy at predicting iteration decisions of oracle labels, as shown in Figure 6.
Adding Iteration Depth. We train a 1.7B TaH with maximum three iteration (TaH-3). TaH-3 yields 5.8% average gain over Standard, and 0.8% over TaH-2. Detailed results are in Appendix A.2.1.
@@ 5.3 Design Choice Exploration
We demonstrate the effectiveness and robustness of TaH by finetuning our model with alternative model architectures and training schemes, or altering the continuation thresholds.
Table 2: Ablation study on architecture designs.
LoRA Residual Attention MATH500 AMC23 Olympiadbench Average
Duo-causal 51.2 32.5 23.9 35.9
Duo-causal 51.6 29.7 22.4 34.6 //-1.3
49.2 22.5 21.2 31.0 //-4.9
Causal 47.8 24.4 19.4 30.5 //-5.4
Model Architecture. (1) LoRA and residual connections. As shown in Table 2, gradually removing LoRA and residual connections leads to consistent performance drops, validating their importance for dynamic iteration. (2) Duo-causal attention. Replacing duo-causal attention with regular causal attention (always attending to depth-1 for full contexts) also degrades performance, confirming the value of allowing tokens to attend to current and all shallower iterations of previous positions.
Training Scheme. (1) Supervision type. Inspired by early exit methods, a common alternative supervises all iteration depths with next-token labels to enable flexible early termination. It enforces accurate prediction at depth 1 even for hard tokens. As shown in Table 3, such token+latent supervision underperforms our token-only approach that supervises only at oracle-determined depths. It aligns with our intuition that different iterations should focus on tokens of different difficulties. (2) Iteration behavior We compare our static oracle strategy $\pi$ with two alternatives. The iter. decider trains the decider first then uses it during backbone training, but suffers from the coupling challenge discussed in Section 4.3. The dynamic recalculates the oracle using the evolving backbone in Equation 7, encountering the same coupling challenge and causing training collapse. These results support our backbone training recipe: using next-token supervision with oracle iteration policy.
Table 3: Ablation study on training schemes.
Supervision Iter. Behavior MATH500 AMC23 Olympiadbench Average
Token-only Oracle 51.2 32.5 23.9 35.9
Token+latent Oracle 49.4 29.6 15.9 31.6 //-4.3
Token-only Iter. decider 44.8 24.1 17.3 28.7 //-7.2
Dynamic 11.0 2.8 2.7 5.5//-30.4
Training Inference Accuracy
Standard Standard 52
AlwaysThink AlwaysThink 38//-14
AlwaysThink TaH-Oracle 77/+/+25
TaH-Oracle TaH-Decider 54/+/+02
TaH-Oracle TaH-Oracle 80/+/+28
Table 4: Impact of iteration strategies on Qwen3-0.6B (first 100 MATH500 samples).
Figure 5: Next-token prediction changes across iterations. Top2 tokens that think-twice most are visualized.
Continuation Threshold. As shown in Figure 4, TaH maintains robust performance across different continuation thresholds and iteration ratios. We empirically set $c_{\text{threshold}}=0.9$ for all evaluations.
@@ 5.4 Behavior Analysis
Latent Overthinking. To analyze latent thinking patterns, we verbalize tokens from all iteration depths using their last-layer hidden states. The oracle method uses the oracle policy $\pi$ from Section 4.3 for iteration decision. (1) Generation. Since ground-truth tokens are unavailable during generation, we use predictions from the stronger DeepSeek-R1-Distill-Qwen-32B model (guo2025deepseek-r1) as proxy labels. Table 4 shows that the oracle policy substantially improves performance by verbalizing correct predictions immediately while iterating only on incorrect ones. With our trained iteration decider approximating the oracle, TaH outperforms both Standard and AlwaysThink baselines. However, the ideal oracle policy achieves even higher gains, indicating future potential. (2) Next-token prediction. We evaluate next-token prediction accuracy on the Open-R1 dataset, using the test model itself as the reference model in $\pi$. Figure 1 reveals that AlwaysThink produces more incorrect than correct revisions, demonstrating latent overthinking. In contrast, oracle-controlled iterations substantially increase correct revisions by selectively targeting hard tokens.
Token Alternation Patterns. We analyze which tokens TaH selects for deeper iteration. On the validation set, But and So emerge as top candidates, with iteration probabilities of 34% and 18%, respectively. These tokens signal critical contrasting or causal relationships, confirming that models benefit from additional processing at logically complex junctures. Figure 5 illustrates how TaH alternates predictions after iteration at these key tokens, suggesting logic refinement behavior. See Appendix A.1.3 for detailed analysis.
Attention Pattern. We visualize the attention pattern of TaH. As discussed in Figure 7 and Appendix A.2.5, the duo-causal attention automatically focuses on different iterations in different heads, extracting broader contexts from multiple depths.
Iteration and FLOPs. Table 9 reports per-token FLOPs and average iteration count. TaH matches the FLOPs of Standard baseline (with average 1.06 iterations per token), far below AlwaysThink (with 2.00 iterations per token) which incurs $2.9\times$ to $4.0\times$ the FLOPs.
@@ 6 Conclusion
We present TaH, a selective latent thinking method that iterates deeper only on hard tokens. Architecturally, TaH introduces duo-causal attention, depth-specific LoRA, and a neural iteration decider to facilitate dynamic depths. An oracle policy guides the stable two-stage training for the tightly coupled LLM backbone and decider. Across five reasoning benchmarks, TaH improves accuracy by 4.0-5.4% over strong baselines with minimal overhead (<<3% additional parameters and $\approx$6% extra iterations), establishing a new paradigm for better reasoning within the current parameter budgets.

### v2

@@ 5.2 Depth-adaptive Model Architecture
During inference, token $i$ verbalizes when $\hat{c}_{i}^{(d)}$ falls below threshold $c_{\text{threshold}}$ or reaches maximum depth $d_{\max}$.
@@ 5.3 Training Scheme
We adopt a two-stage training scheme. We first fine-tune the backbone model to support selective latent iteration, then train the iteration decider. Both stages are aligned to the same oracle iteration policy.
Motivation. As shown in Figure 2(c), the backbone LLM and the iteration decider are tightly coupled, making joint training unstable. Specifically, the backbone’s prediction quality across iterations determines the optimal depth, while the decider controls the backbone’s KV cache and output depth. To stabilize training, we train the two components sequentially under a fixed oracle policy $\pi$ (validated in Section 4).
Oracle Iteration Policy $\pi$.
Our goal is to iterate only on hard tokens that a standard supervised fine-tuned (SFT) model would mispredict on the first pass. Thus, we define $\pi$ with SFT model as the reference: we trigger an additional iteration when its top-1 next-token prediction differs from the ground-truth token. In principle, one could instead define $\pi$ using the current looped model itself (i.e., an on-policy oracle), but we find this to be empirically unstable (Section 6.3).
@@
We freeze the backbone model $(\theta,\Delta)$ and train the iteration decider $\phi$ to imitate the oracle policy’s continuation decisions. We minimize weighted binary cross-entropy:
$\mathcal{L}_{\text{dec}}(\phi)\;=\;-\sum_{i,d}w_{d}\Big[c_{i}^{(d)}\log\hat{c}_{i}^{(d)}+(1-c_{i}^{(d)})\log(1-\hat{c}_{i}^{(d)})\Big],$
where the sum ranges over tokens $i$ and depths $d=1,\ldots,\min\{d_{\max}-1,d_{i}^{\pi}\}$. Here $c_{i}^{(d)}$ is the ground-truth continuation label, $\hat{c}_{i}^{(d)}$ is the predicted probability, and $w_{d}$ is the class weight for label imbalance (ratio of stop to continue labels).
The two-stage scheme stabilizes training by decoupling backbone learning (conditioned on a fixed $\pi$) from iteration policy learning (imitation of $\pi$).
@@ 6.1 Setup
We present key experiment configurations here, with detailed setups in Appendix A.1.
Baselines. We compare diverse methods under equal parameter budgets, using Qwen3-{0.6B,1.7B,4B}-Base (Yang et al., 2025) as backbones. We compare TaH over the following baselines: (1) Standard, which always verbalizes directly and reduces to the standard Qwen model; (2) SoftThink, a latent optimization method, implemented following official design (Zhang et al., 2025) on top of the Standard model; (3) AlwaysThink, similar to TaH, but always iterates twice for all tokens during training and inference; (4) Ouro, a looped transformer that can scale iteration depths, implemented following official design (Zhu et al., 2025). Unless otherwise specified, TaH, Ouro, and AlwaysThink all use a maximum of two iterations, fine-tuned from the same Qwen3 backbone on the same training data and recipe. Ouro reaches its highest performance when set to maximum iterations, so we report results under this setup.
TaH Setup. To match the total parameter count of TaH (with LoRA and decider) with that of baselines, we prune one layer from the LLM backbone before training. The layer is chosen to minimize the increase in validation loss. We also report results for an unpruned variant, TaH+, which adds less than 3% extra parameters from LoRA and decider. The detailed parameter composition is shown in Table 4. Following (Fu et al., 2025), we set the continuation threshold $c_{\text{threshold}}=0.9$ with about 7% of tokens being iterated twice. The oracle policy $\pi$ uses Qwen3-0.6B, 1.7B, and 4B as reference models, respectively.
Training Scheme. All models are trained on the balanced Open-R1 (Hugging Face, 2025) mixture (math, QA, and code; 100K samples) using supervised fine-tuning. To fit memory and compute limits, we exclude responses longer than 8,192 tokens; 4B models additionally truncate at 4,096 tokens; all other training settings follow the official Open-R1 script. Each method is sufficiently trained for 5 epochs, and we select the checkpoint with the lowest validation loss as the final model. All backbones are initialized from the corresponding Qwen3-Base model.
Evaluation Setup. We evaluate across challenging reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021b), AMC23 (American Mathematics Competitions), AIME25 (American Invitational Mathematics Examination), OlympiadBench (He et al., 2024), MBPP++ (Austin et al., 2021), HumanEval++ (Chen et al., 2021), GPQA-Diamond (denoted as GPQA) (Rein et al., 2023), and MMLU-STEM  (Hendrycks et al., 2021a). The maximum generation length is set to 8,192 tokens for all benchmarks, except GSM8K and MMLU-STEM which use 4,096 due to their simpler problems and larger size. Performance is reported as pass@1 under a zero-shot CoT setting, using sampling temperature 0.6. We generate one sample per problem for large datasets (MATH500, OlympiadBench etc.), and eight samples per problem for small datasets (AMC23, AIME25).
@@ 6.2 Performance
Benchmark Evaluation. We validate TaH’s reasoning ability across all nine benchmarks. Table 5 presents performance results for models at 0.6B, 1.7B, and 4B parameter sizes. Compared with the strong Standard Qwen3 baselines, we observe that existing approaches (AlwaysThink, SoftThink, and Ouro) yield only marginal improvements when fine-tuned from base. In contrast, TaH achieves consistent gains across benchmarks. For 0.6B and 1.7B models, TaH delivers average improvements of 3.0% and 3.8% over Standard, respectively; TaH+, which adds less than 3% additional parameters, further pushes these gains to 5.3% and 6.2%. Compared to concurrent work Ouro, TaH and TaH+ achieve 3.8–4.4% and 6.1–6.8% gains, respectively. For 4B models, which are trained with 4K context length due to resource limits, TaH and TaH+ also achieve gains of 1.7% and 2.2%, with potential for further improvement at longer training contexts.
Table 2: Accuracy comparison across benchmarks. Best results are highlighted in bold. 4B models trained with $\leq$4K lengths due to resource constraints; AlwaysThink is excluded due to Out-Of-Memory (OOM) during training.
Standard SoftThink Ouro AlwaysThink TaH TaH+
0.6B
AIME25 1.9 2.9 2.1 1.3 2.1 4.6
OlympiadBench 15.4 14.0 14.2 12.6 19.1 20.6
AMC23 22.7 22.2 19.7 21.9 24.1 24.7
MATH500 39.9 39.6 37.4 37.8 46.2 51.8
GSM8K 58.2 55.9 56.6 52.6 63.6 67.6
GPQA 31.1 24.7 35.4 30.8 29.0 31.3
MMLU-STEM 54.2 53.0 54.0 51.4 56.4 59.0
HumanEval++ 16.8 14.3 18.9 9.1 21.6 22.0
MBPP++ 28.8 29.5 23.5 13.8 33.9 35.1
Average 29.9 28.5 29.1 25.7 32.9/+3.0/+3.0 35.2/+5.3/+5.3
1.7B
AIME25 10.8 5.4 10.8 7.5 13.8 15.4
OlympiadBench 33.8 30.7 31.3 31.0 37.2 37.6
AMC23 39.7 40.3 40.6 40.9 40.9 48.4
MATH500 67.8 64.8 68.2 63.2 71.4 72.6
GSM8K 80.2 80.0 79.8 74.2 84.8 84.5
GPQA 30.3 33.3 32.8 30.5 33.3 39.4
MMLU-STEM 74.1 73.5 72.4 69.6 74.8 76.6
HumanEval++ 39.0 43.3 40.9 16.4 50.0 51.5
MBPP++ 51.9 49.1 45.5 25.6 55.3 57.5
Average 47.5 46.7 46.9 39.9 51.3/+3.8/+3.8 53.7/+6.2/+6.2
4B
AIME25 24.2 25.8 25.0 27.1 27.9
OlympiadBench 50.4 50.2 51.4 50.5 52.6
AMC23 64.2 63.1 64.4 69.7 68.1
MATH500 84.2 85.0 83.8 85.8 85.6
GSM8K 91.7 92.5 90.7 91.0 91.7
GPQA 46.2 50.3 50.0 48.5 49.0
MMLU-STEM 85.4 86.0 85.9 86.2 86.6
HumanEval++ 69.2 69.2 70.7 70.1 72.0
MBPP++ 65.9 66.7 66.7 67.7 68.1
Average 64.6 65.4 65.4 66.3/+1.7/+1.7 66.8/+2.2/+2.2
Training Dynamics. During stage 1 (LLM backbone training), TaH performs iterations according to the oracle policy. As shown in Figure 4(a), it converges notably faster than the Standard baseline and also achieves much lower validation perplexity. During stage 2 (iteration-decider training), the neural decider successfully imitates the oracle strategy. It reaches about 83% accuracy at predicting the oracle’s iteration decisions, as shown in Figure 9.
@@ 6.3 Design Choice Exploration
We demonstrate the effectiveness and robustness of each design choice of TaH through ablation studies. All experiments in this section train TaH and its variants on the math subset of Open-R1 and evaluate on MATH500, AMC23, and OlympiadBench.
Model Architecture. (1) Iteration Scheme. As shown in Table 13, decider-based iteration outperforms the Always-1 and Always-2 alternatives, confirming the practical benefits of selective iteration, even with imperfect decisions from neural decider. Note that for Always-1, duo-causal attention degenerates to regular causal attention. (2) Duo-Causal Attention. Replacing duo-causal attention with standard causal attention variants causes significant drops: (a) attending only to the first iteration (Causal@iter1) drops by 5.4%; (b) attending only to the current iteration (Causal@current) drops even more, by 8.5%. They confirm duo-causal attention’s essential role in cross-depth information flow. (3) Depth Adapters. Removing LoRA and residual connections leads to consistent drops, confirming their beneficial roles in objective transition across iterations.
Training Scheme. (1) Supervision type. Some previous work supervises all iteration depths with next-token labels, while TaH only supervises final output tokens. As shown in Table 13, such token+latent supervision underperforms TaH. This aligns with our intuition that different iterations should focus on their respective objectives. (2) Iteration policy during LLM training. We compare our static oracle strategy $\pi$ with two alternatives. The decider-based approach trains the iteration decider first, then uses it during backbone training. It suffers from the coupling challenge discussed in Section 5.3. The dynamic approach recalculates the oracle using the evolving backbone in Equation 5, facing the same coupling challenge and causing training collapse. These results support our selection of the static oracle policy. (3) Discrepancy metric in $\pi$. We compare three discrepancy metrics to trigger latent iteration in $\pi$: top-1 mismatch, entropy, and cross-entropy (detailed definitions in Appendix A.3.1). As Table 13 shows, top-1 mismatch yields the best result, confirming its empirical effectiveness.
Continuation Threshold. As shown in Figure 4(b), TaH maintains robust performance across different continuation thresholds and iteration ratios. We empirically set $c_{\text{threshold}}=0.9$ for all evaluations.
Iteration Depth. We train a 1.7B TaH-3 with a maximum of three iterations. TaH-3 yields 5.8% average gain over Standard, and 0.8% over TaH-2, as detailed in Appendix A.2.5.
@@ 6.4 Behavior Analysis
Token Alternation Patterns. We analyze which tokens TaH selects for deeper iteration. On the validation set, But and So emerge as top candidates, with iteration probabilities of 34% and 18%, respectively. These tokens signal critical contrasting or causal relationships, confirming that models benefit from additional processing at logically complex junctures. Figure 3(a) illustrates how TaH alternates predictions after iteration at these key tokens, suggesting logic refinement behavior. See Appendix A.3.4 for detailed analysis.
Attention Pattern. We visualize the attention pattern of TaH in Figure 10 and Appendix A.3.6. The duo-causal attention learns to focus on different iterations in different heads, extracting broader contexts from multiple depths.
(a) Next-token prediction changes across iterations. Top-2 tokens that think-twice most frequently are visualized.
(b) Distribution of token prediction accuracy across iterations for Ouro and TaH.
Figure 3: Behavior analysis visualizations.
Accuracy Landscape. Figure 3 shows NTP accuracy across iterations. Because Ouro trains all iterations to predict all tokens, predictable tokens across depths largely overlap, leaving more tokens unpredictable by any iteration. TaH instead specializes deeper iterations for hard tokens, reducing overlap and improving coverage under selective iteration.
@@ 7 Conclusion
We present TaH, a selective latent iteration method that simultaneously improves model performance and efficiency. Architecturally, TaH introduces duo-causal attention, depth-specific LoRA, and a neural iteration decider, enabling effective selective iteration. An oracle policy guides the stable two-stage training of the tightly coupled LLM backbone and decider. Across nine benchmarks, TaH improves accuracy by 5.3–6.2% over strong baselines with minimal overhead, establishing a new direction for improved reasoning under parameter constraints.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 62506197, 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164, 92364201), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).